# NetMamba: Efficient Network Traffic Classification via Pre-training Unidirectional Mamba

### v1

@@ Abstract
Network traffic classification is a crucial research area aiming to enhance service quality, streamline network management, and bolster cybersecurity. To address the growing complexity of transmission encryption techniques, various machine learning and deep learning methods have been proposed. However, existing approaches encounter two main challenges. Firstly, they struggle with model inefficiency due to the quadratic complexity of the widely used Transformer architecture. Secondly, they suffer from unreliable traffic representation because of discarding important byte information while retaining unwanted biases. To address these challenges, we propose NetMamba, an efficient linear-time state space model equipped with a comprehensive traffic representation scheme. We replace the Transformer with our specially selected and improved Mamba architecture for the networking field to address efficiency issues. In addition, we design a scheme for traffic representation, which is used to extract valid information from massive traffic while removing biased information. Evaluation experiments on six public datasets encompassing three main classification tasks showcase NetMamba’s superior classification performance compared to state-of-the-art baselines. It achieves up to 4.83% higher accuracy and 4.64% higher f1 score on encrypted traffic classification tasks. Additionally, NetMamba demonstrates excellent efficiency, improving inference speed by 2.24 times while maintaining comparably low memory usage. Furthermore, NetMamba exhibits superior few-shot learning abilities, achieving better classification performance with fewer labeled data. To the best of our knowledge, NetMamba is the first model to tailor the Mamba architecture for networking.
@@ I Introduction
Recently, pre-training has emerged as a prevalent model training paradigm in natural language processing (NLP) [7] and computer vision (CV) [8]. Motivated by this trend, several Transformer-based pre-trained traffic models [9, 10, 11] have been developed to learn generic traffic representations from extensive unlabeled data and then fine-tune for specific downstream tasks using limited labeled traffic data. However, these existing models face two significant challenges: 1) Limited Model Efficiency: the widely used Transformer architecture employs a quadratic self-attention mechanism, leading to substantial computational and memory costs on long sequences; 2) Unreliable Traffic Representation: current methodologies cannot adequately and accurately represent raw traffic data due to discarding crucial byte information and preserving unwanted biases.
@@
To improve model efficiency, we replace the Transformer with a different backbone architecture, namely Mamba. Mamba, a liner-time state space model for sequence modeling, has achieved notable success across various domains, including natural language processing [12], computer vision [13] and graph understanding [14]. This suggests a promising potential for applying Mamba to the network domain. However, adapting Mamba for efficient and robust network traffic analysis requires the selection of an appropriate architecture from the existing heterogeneous Mamba variants. By carefully testing different variants of Mamba, we found that the original unidirectional Mamba, without omnidirectional scans or redundant blocks, is well-suited for learning latent patterns within sequential network traffic. To further enhance the model’s performance and robustness, we incorporate positional embeddings and pre-training, specially designed for networking.
@@
To enhance traffic representation, we design a more comprehensive and reliable scheme. This scheme retains valuable packet content within both headers and payloads while eliminating unwanted biases through various methods, including packet anonymizing, byte balancing and stride cutting, thereby improving traffic classification capabilities.
@@
Based on the aforementioned designs, we propose NetMamba, a novel pre-trained unidirectional state space model for efficient and robust network traffic classification. Specifically, NetMamba initially extracts hierarchical flow information from raw traffic and converts it into a stride sequence, which serves as the model’s input. Subsequently, NetMamba undergoes self-supervised pre-training on large unlabeled datasets using a masked autoencoder structure, which is designed to learn generic representations of traffic data through reconstructing masked strides. Finally, the decoder is replaced with a multi-layer perceptron head, and NetMamba is fine-tuned on limited labeled data to refine traffic representations and adapt to downstream traffic classification tasks. Extensive experiments conducted on publicly available datasets demonstrate the effectiveness and efficiency of NetMamba. In all classification tasks, NetMamba achieves an accuracy rate of nearly 99% (some over 99%), and, compared to existing baselines, it improves inference speed by 2.24 times while maintaining low GPU memory usage. Furthermore, NetMamba exhibits superior few-shot learning capabilities in comparison to other pre-training models, achieving better performance with fewer labeled data.
@@
    We conduct extensive experiments across a range of traffic classification tasks, including encrypted application classification, encrypted malware classification, and attack traffic classification. An overall comparison, along with detailed evaluations—encompassing ablation studies, efficiency analyses, and few-shot learning investigations—is provided. These insights could illuminate paths for future research. Additionally, The code of NetMamba is publicly available  [^1] https://github.com/wangtz19/NetMamba.
@@ II-A Transformer-based Traffic Classification
To leverage unlabeled data effectively, several pre-training models have been proposed. Inspired by BERT’s pre-training methodology in natural language processing, PERT[17] and ET-BERT[9] process raw traffic bytes using tokenization, apply masked language modeling to learn traffic representations, and fine-tune the models for downstream tasks. Similarly, YaTC[10] and FlowMAE[18] adopt the widely-used MAE pre-training approach from computer vision, which involves patch splitting for byte matrices, capturing traffic correlations through masked patch reconstruction, and subsequent fine-tuning.
@@ II-B Mamba-based Representation Learning
III Motivation
III-A Improving Model Efficiency
State-of-the-art methods in traffic analysis primarily use the Transformer architecture, which employs self-attention to calculate correlations within a sequence. While the Transformer is effective for sequence modeling, its quadratic computational and memory complexity relative to sequence length leads to significant time and memory costs with longer sequences [13, 26]. Given the requirement for real-time online traffic classification and the limited resources of typical network devices, there is an urgent need for a computationally efficient and memory-optimized traffic model.
Mamba, a novel sequence model, offers linear time complexity and optimized memory costs while maintaining strong modeling capabilities. Motivated by successful adaptations of domain-specific Mamba variants such as Vim[13], GraphMamba[14], and VL-Mamba[25], we propose NetMamba as a more efficient solution for traffic classification.
III-B Enhancing Traffic Representation
In real-world scenarios, massive raw network traffic encompasses a wide range of data categories that differ in upper applications, carried protocols, or transmission purposes (such as benign or malicious). Therefore, it is crucial to devise a robust representation scheme with appropriate granularity to enhance traffic understanding and monitoring capabilities.
Traditional machine learning methods are frequently constrained by restricted model parameters and limited fitting capabilities. Consequently, they commonly resort to utilizing compressed statistical features at the packet or flow level, such as distributions of packet sizes or inter-arrival times. However, these features often suffer from excessive compression, resulting in the loss of vital information inherent in raw datagrams. This can lead to model degradation or even invalidation, especially in complex traffic scenarios.
Recent advancements in deep learning have endeavored to leverage raw traffic bytes. As depicted in Table I, these methods face limitations, such as neglecting IP headers, preserving non-anonymous biases, overlooking byte balance, or employing improper data-splitting techniques.
To address these issues, we propose a novel network traffic representation scheme. This approach retains header information, mitigates IP biases, ensures balanced byte allocation, and tailors data-splitting techniques. Ultimately, this method preserves hierarchical traffic information while eliminating biases effectively.
@@ IV Preliminaries
IV-A State Space Models
@@ IV-B Discretization
where $\mathbf{\overline{A}}=\exp(\Delta\mathbf{A})$ and $\mathbf{\overline{B}}=(\Delta\mathbf{A})^{-1}(\exp(\Delta\mathbf{A})-\mathbf{I})\cdot\Delta\mathbf{B}\approx(\Delta\mathbf{A})^{-1}(\Delta\mathbf{A})\Delta\mathbf{B}=\Delta\mathbf{B}$ represent the discretized parameters, with $\Delta$ denoting the discretization step size. This recurrent formulation, known for its linear time complexity, is suitable for model inference but lacks parallelizability during training.
By expanding Equation 2, SSMs can be transformed into convolutional formulations as follows:
$\mathbf{\overline{K}}$ $=(\mathbf{C}\mathbf{\overline{B}},\mathbf{C}\mathbf{\overline{A}}\mathbf{\overline{B}},\dots,\mathbf{C}\mathbf{\overline{A}}^{L-1}\mathbf{\overline{B}})$ (3)
$y$ $=x*\mathbf{\overline{K}}$
where $L$ represents the length of the input sequence $x$, * denotes the convolution operation, and $\mathbf{\overline{K}}\in\mathbb{R}^{L}$ refers to a structured convolutional kernel. This convolutional representation solves the computational parallelization dilemma encountered in the recurrent version.
IV-C Selection Mechanism
@@ V NetMamba Framework
This section overviews the architecture of NetMamba (see Figure 1), providing a comprehensive blueprint for the detailed design presented in § VI. Initially, NetMamba extracts hierarchical information from raw binary traffic and converts it into stride-based representation. Inspired by the Masked AutoEncoders (MAE) pre-training model in computer vision, NetMamba employs a dual-stage training approach. Specifically, self-supervised pre-training is utilized to acquire traffic representation, while supervised fine-tuning is employed to tailor the model for downstream traffic understanding tasks.
V-1 Traffic Representation Phase
@@
To enhance domain knowledge within networks, NetMamba adopts a stride-based methodology to represent network traffic comprehensively. Initially, network traffic is segmented into distinct flows, categorized by their 5-tuple attributes: Source IP, Destination IP, Source Port, Destination Port, and Protocol. Subsequently, for each packet within a flow, fixed-sized segments comprising header and payload bytes are extracted. To collect more comprehensive traffic information without compromising model efficiency due to excessively long packet sequences, we follow approaches outlined in prior studies[9, 10], which involve selectively utilizing specific packets within a flow. Specifically, we aggregate the bytes from certain initial packets of each flow into a unified byte array. This array integrates information across byte, packet, and flow levels, offering a holistic perspective on traffic characteristics.
@@
This byte array forms the foundation for segmenting non-overlapping flow strides. It preserves semantic relationships between adjacent bytes, effectively mitigating biases introduced by conventional patch-splitting methods, as well as addressing out-of-vocabulary issues commonly associated with tokenization processes. Further design intricacies regarding traffic representation are elucidated in § VI-A.
V-2 Pre-training Phase
@@
To acquire generic encodings of network domain knowledge based on flow stride representations, NetMamba undergoes pre-training using extensive unlabeled network traffic data. Specifically, NetMamba employs a masked autoencoder (MAE) architecture, incorporating multiple unidirectional Mamba blocks in both the traffic encoder and decoder. The internal workings of the Mamba block are detailed in § VI-B2. Following the original design of MAE, NetMamba is characterized by an asymmetric structure, comprising a heavyweight encoder and a lightweight decoder. This configuration is chosen to cultivate a more efficient traffic representation encoder.
@@
During the pre-training phase, flow strides are processed through several sequential steps. Initially, they are concatenated with a trailing class token and then mapped into stride embeddings. This is followed by the addition of positional embeddings and the random masking of strides before their introduction to the NetMamba encoder. The encoder focuses solely on visible strides, grasping inherent relationships and generating an output traffic representation. Conversely, the decoder endeavors to reconstruct the masked strides using the output from the encoder in conjunction with dummy masked tokens. The optimization of the pre-training phase is achieved by minimizing the reconstruction loss, particularly for the masked strides, thereby ensuring that the model captures robust representations of traffic patterns. Detailed insights into the pre-training strategy are provided in § VI-C.
V-3 Fine-tuning Phase
@@
For accurately capturing traffic patterns and understanding task requirements specific to a downstream task, NetMamba undergoes fine-tuning using labeled traffic data. During the fine-tuning phase, the decoder of NetMamba is replaced by a multi-layer perceptron (MLP) head to facilitate classification tasks. With the removal of the reconstruction task, all embedded flow strides become visible to the encoder. As the unidirectional Mamba block processes sequence information in a front-to-back manner, the trailing class token, after being processed by the encoder, aggregates the overall traffic characteristics. Subsequently, NetMamba forwards only this class token to the MLP-based classifier.
@@
Following pre-training, the encoder of NetMamba demonstrates significant adaptability when fine-tuned with a limited quantity of labeled, task-specific traffic data. This adaptability allows NetMamba to efficiently transition to a variety of downstream tasks, including but not limited to application classification and attack detection. For additional information concerning the fine-tuning process, please refer to Section VI-D.
VI Detailed Design
This section details the design of NetMamba, covering the traffic representation scheme, the NetMamba model architecture, and the pre-training and fine-tuning strategies. Key hyper-parameters of NetMamba are summarized in Table II.
@@ VI Detailed Design
VI-A Traffic Representation
VI-A1 Flow Splitting
@@
VI-A2 Packet Parsing
For each flow, all packets are processed through several sequential operations to preserve valuable information and eliminate unnecessary interference. When narrowing down the scope for analyzing traffic data related to specific applications or services, we exclude all packets carried by non-IP protocols, such as Address Resolution Protocol (ARP) and Dynamic Host Configuration Protocol (DHCP). Considering the critical information contained within both the IP header (e.g., the total length field) and the payload (text content for upper-level protocols), we choose to retain these elements. Furthermore, to mitigate biases introduced by identifiable information, all packets are anonymized through the removal of potential Ethernet headers and the masking of IP addresses.
VI-A3 Packet Cropping & Padding, and Concatenating
Given the variability in packet size within the same flow and the fluctuation in both header length (including the IP header and any potential upper-layer headers) and payload length within individual packets, problematic scenarios often arise. For instance, the first long packet can occupy the entire limited model input array, or excessively long payloads can dominate the byte information within shorter headers. Therefore, it is essential to standardize packet sizes by assigning uniform sizes to all packets and fixed lengths to both packet headers and payloads. Specifically, we select the first $M$ packets from a single flow, setting the header length to $N_{h}$ bytes and the payload length to $N_{p}$ bytes. Any packet exceeding this length will be cropped, while shorter packets will be padded to meet these specifications.
@@ VI-A3 Packet Cropping & Padding, and Concatenating
VI-A4 Stride Cutting
Given the significant computational and memory demands posed by a byte array with $L_{b}$ (typically greater than 1000) elements, it becomes imperative to explore further compression techniques to enhance the efficiency of model training and inference. Traditional methods often involve reshaping the byte array into a square matrix and employing two-dimensional patch splitting, a practice borrowed from computer vision. However, this approach inadvertently introduces biases by grouping vertically adjacent bytes that are semantically unrelated.

### v3

@@ Abstract
Network traffic classification is a crucial research area aiming to enhance service quality, streamline network management, and bolster cybersecurity. To address the growing complexity of transmission encryption techniques, various machine learning and deep learning methods have been proposed. However, existing approaches face two main challenges. Firstly, they struggle with model inefficiency due to the quadratic complexity of the widely used Transformer architecture. Secondly, they suffer from inadequate traffic representation because of discarding important byte information while retaining unwanted biases. To address these challenges, we propose NetMamba, an efficient linear-time state space model equipped with a comprehensive traffic representation scheme. We adopt a specially selected and improved unidirectional Mamba architecture for the networking field, instead of the Transformer, to address efficiency issues. In addition, we design a traffic representation scheme to extract valid information from massive traffic data while removing biased information. Evaluation experiments on six public datasets encompassing three main classification tasks showcase NetMamba’s superior classification performance compared to state-of-the-art baselines. It achieves an accuracy rate of nearly 99% (some over 99%) in all tasks. Additionally, NetMamba demonstrates excellent efficiency, improving inference speed by up to 60 times while maintaining comparably low memory usage. Furthermore, NetMamba exhibits superior few-shot learning abilities, achieving better classification performance with fewer labeled data. To the best of our knowledge, NetMamba is the first model to tailor the Mamba architecture for networking.
@@ I Introduction
Recently, pre-training has emerged as a prevalent model training paradigm in natural language processing (NLP) [7] and computer vision (CV) [8]. Motivated by this trend, several Transformer-based pre-trained traffic models [9, 10, 11] have been developed to learn generic traffic representations from extensive unlabeled data and then fine-tune for specific downstream tasks using limited labeled traffic data. However, these existing models face two significant challenges: 1) Limited Model Efficiency: state-of-the-art methods in traffic analysis primarily use Transformer architecture, which employs a quadratic self-attention mechanism to calculate correlations within a sequence. This leads to substantial computational and memory costs on long sequences[12, 13]. Consequently, these models are unsuitable for real-time online traffic classification and cannot operate efficiently with the limited resources of typical network devices. 2) Inadequate Traffic Representation: current methodologies fail to adequately and accurately represent raw traffic data due to discarding crucial byte information and preserving unwanted biases. As a result, these unreliable schemes impair classification performance or even cause model failure in complex traffic scenarios.
To address these challenges, we propose NetMamba, an efficient linear-time state space model equipped with a comprehensive traffic representation scheme, aiming to accurately perform network traffic classification tasks with higher inference speed and lower memory usage.
@@
To improve model efficiency, we use the Mamba architecture for the model backbone instead of the Transformer. Mamba[14], a liner-time state space model for sequence modeling, has achieved notable success across various domains, including natural language processing [15], computer vision [12] and graph understanding [16]. This suggests promising potential for applying Mamba to the network domain. However, adapting Mamba for efficient and robust network traffic analysis requires selecting the appropriate architecture from the existing heterogeneous and complex Mamba variants. By carefully testing different variants of Mamba, we found that the original unidirectional Mamba[14], without omnidirectional scans or redundant blocks, is well-suited for efficiently learning latent patterns within sequential network traffic. To further enhance the model’s performance and robustness, we incorporate positional embeddings and pre-training strategies specially designed for networking.
@@
To enhance traffic representation, we design a more comprehensive and reliable scheme. This scheme retains valuable packet content within both headers and payloads while eliminating unwanted biases through various methods, including packet anonymizing, byte allocation balancing and stride-based data cutting, thereby improving traffic classification capabilities.
@@
Specifically, NetMamba initially extracts hierarchical flow information from raw traffic and converts it into a stride sequence, which serves as the model’s input. Subsequently, NetMamba undergoes self-supervised pre-training on large unlabeled datasets using a masked autoencoder structure, which is designed to learn generic representations of traffic data through reconstructing masked strides. Finally, the decoder is replaced with a multi-layer perceptron head, and NetMamba is fine-tuned on limited labeled data to refine traffic representations and adapt to downstream traffic classification tasks. Extensive experiments conducted on publicly available datasets demonstrate the effectiveness and efficiency of NetMamba. In all classification tasks, NetMamba achieves an accuracy rate of nearly 99% (some over 99%), and, compared to existing baselines, it improves inference speed by up to 60 times while maintaining low GPU memory usage. Furthermore, NetMamba exhibits superior few-shot learning capabilities in comparison to other pre-training models, achieving better performance with fewer labeled data.
@@
    We conduct extensive experiments across a range of traffic classification tasks. An overall comparison, along with detailed evaluations—encompassing ablation studies, efficiency analyses, and few-shot learning investigations—is provided. These insights could illuminate paths for future research. Additionally, the code of NetMamba is publicly available  [^1] https://github.com/wangtz19/NetMamba.
@@ II-A Transformer-based Traffic Classification
To leverage unlabeled data effectively, several pre-trained models have been proposed. Inspired by BERT’s pre-training methodology in natural language processing, PERT [19] and ET-BERT [9] process raw traffic bytes using tokenization, apply masked language modeling to learn traffic representations, and fine-tune the models for downstream tasks. Similarly, YaTC [10] and FlowMAE [20] adopt the widely-used MAE pre-training approach from computer vision, which involves patch splitting for byte matrices, capturing traffic correlations through masked patch reconstruction, and subsequent fine-tuning.
@@ II-B Mamba-based Representation Learning
II-C Traffic Representation Schemes
In real-world scenarios, massive raw network traffic encompasses a wide range of data categories that vary in upper applications, carried protocols, or transmission purposes. Therefore, a robust representation scheme with appropriate granularity is crucial for accurate traffic understanding.
Traditional machine learning methods [1, 2, 3, 27, 28], constrained by limited model parameters and fitting capabilities, commonly resort to utilizing compressed statistical features at the packet or flow level, such as distributions of packet sizes or inter-arrival times. However, these features often suffer from excessive compression, resulting in the loss of vital information inherent in raw datagrams.
Recent advancements in deep learning have endeavored to utilize raw traffic bytes. However, as shown in Table I, these methods face limitations. They often neglect crucial information in packet headers and introduce unwanted biases by retaining biased IP addresses, ignoring byte balance, or using improper data-splitting techniques.
To address these issues, we propose a novel network traffic representation scheme. Our approach remedies the aforementioned shortcomings, preserving hierarchical traffic information while effectively eliminating biases.
@@ III Preliminaries
III-A State Space Models
@@ III-B Discretization
where $\mathbf{\overline{A}}=\exp(\Delta\mathbf{A})$ and $\mathbf{\overline{B}}\approx\Delta\mathbf{B}$ represent the discretized parameters, with $\Delta$ denoting the discretization step size. This recurrent formulation is characterized by linear time complexity.
III-C Selection Mechanism
@@ IV NetMamba Framework
This section overviews the framework of NetMamba (see Figure 1), providing a comprehensive blueprint for the detailed design presented in § V and § VI. Initially, NetMamba extracts hierarchical information from raw binary traffic and converts it into stride-based representation. Inspired by the Masked AutoEncoders (MAE) pre-training model in computer vision, NetMamba employs a dual-stage training approach. Specifically, self-supervised pre-training is utilized to acquire traffic representation, while supervised fine-tuning is employed to tailor the model for downstream traffic understanding tasks.
IV-1 Traffic Representation Phase
@@
To enhance domain knowledge within networks, NetMamba adopts a stride-based methodology to represent key content within network traffic. Initially, traffic data is segmented into distinct flows, categorized by their 5-tuple attributes: Source IP, Destination IP, Source Port, Destination Port, and Protocol. Fixed-sized segments of header and payload bytes are then extracted for each packet within a flow. To collect more comprehensive traffic information without compromising model efficiency due to excessively long packet sequences, we follow approaches outlined in prior studies[9, 10], which involve selectively utilizing specific packets within a flow. Specifically, bytes from the initial packets of each flow are aggregated into a unified byte array, integrating information across byte, packet, and flow levels for a comprehensive view of traffic characteristics.
@@
This byte array forms the foundation for segmenting non-overlapping flow strides. It preserves semantic relationships between adjacent bytes, effectively mitigating biases introduced by conventional patch-splitting methods, as well as addressing out-of-vocabulary issues commonly associated with tokenization processes. Further design intricacies regarding traffic representation are elucidated in § V.
IV-2 Pre-training Phase
@@
To acquire generic encodings of network domain knowledge based on flow stride representations, NetMamba undergoes pre-training using extensive unlabeled network traffic data. Specifically, NetMamba utilizes a masked autoencoder (MAE) architecture, incorporating multiple unidirectional Mamba blocks in both its encoder and decoder, as detailed in § VI-A2.
@@
During pre-training, flow strides undergo several sequential steps: concatenation with a trailing class token, mapping into stride embeddings, addition of positional embeddings, and random masking. The encoder focuses solely on visible strides, grasping inherent relationships and generating an output traffic representation. The decoder then reconstructs the masked strides using the encoder’s output and dummy tokens. Pre-training is optimized by minimizing the reconstruction loss for the masked strides, ensuring the model learns robust traffic patterns. Detailed insights into the pre-training strategy are provided in § VI-B.
IV-3 Fine-tuning Phase
@@
For accurately capturing traffic patterns and understanding downstream task requirements, NetMamba undergoes fine-tuning using labeled traffic data. During this phase, the decoder of NetMamba is replaced by a multi-layer perceptron (MLP) head to facilitate classification tasks. With the removal of the reconstruction task, all embedded flow strides become visible to the encoder. As the unidirectional Mamba block processes sequence information in a front-to-back manner, the trailing class token, after being processed by the encoder, aggregates the overall traffic characteristics. Subsequently, NetMamba forwards only this class token to the MLP-based classifier.
@@
Post pre-training, NetMamba’s encoder exhibits significant adaptability when fine-tuned with limited labeled data, enabling efficient transition to various downstream tasks such as application classification and attack detection. Additional details on the fine-tuning process are provided in § VI-C.
V Traffic Representation
This section provides detailed information about the traffic representation scheme used by NetMamba. The key hyper-parameters are listed in Table II.
@@ V Traffic Representation
V-1 Flow Splitting
@@
V-2 Packet Parsing
For each flow, all packets are processed through several sequential operations to preserve valuable information and eliminate unnecessary interference. When narrowing down the scope for analyzing traffic data related to specific applications or services, we exclude all packets carried by non-IP protocols, such as Address Resolution Protocol (ARP) and Dynamic Host Configuration Protocol (DHCP). Considering the critical information contained within both the packet (e.g., the total length field) and the payload (text content for upper-level protocols), we choose to retain these elements. Furthermore, to mitigate biases introduced by identifiable information, all packets are anonymized through the removal of Ethernet headers and the masking of IP addresses.
V-3 Packet Cropping & Padding, and Concatenating
Given the variability in packet size within the same flow and the fluctuation in both header length (including the IP header and any upper-layer headers) and payload length within individual packets, problematic scenarios often arise. For instance, the first long packet can occupy the entire limited model input array, or excessively long payloads can dominate the byte information within shorter headers. Therefore, it is essential to standardize packet sizes by assigning uniform sizes to all packets and fixed lengths to both packet headers and payloads. Specifically, we select the first $M$ packets from a single flow, setting the header length to $N_{h}$ bytes and the payload length to $N_{p}$ bytes. Any packet exceeding this length will be cropped, while shorter packets will be padded to meet these specifications.
@@ V-3 Packet Cropping & Padding, and Concatenating
V-4 Stride Cutting
Given the significant computational and memory demands posed by a byte array with $L_{b}$ (typically greater than 1000) elements, it becomes imperative to explore further compression techniques to enhance the efficiency of model training and inference. Traditional methods often involve reshaping the byte array into a square matrix and employing two-dimensional patch splitting, a practice borrowed from computer vision. However, this technique unintentionally introduces biases by grouping vertically adjacent bytes that are semantically unrelated, as they are not naturally contiguous in the sequential traffic data.